# Coupling between small polarons and ferroelectricity in BaTiO3

### v3

@@ III.1 Polarons : Electrons and Hole Polarons in BaTiO3
Because we now understand the factors that affect the stability of polarons in the material, it is essential to examine the resulting structural distortions introduced by localized charge carriers. Substantial structural distortions were observed due to the formation of polarons and they displayed variations based on the type of polaron and the BaTiO3 polymorph. In cubic BaTiO3, the formation of an electron polaron is accompanied by an equal increase in the Ti-O bond lengths around the polaronic sites. The elongation of the Ti-O bonds around the polaronic site was approximately 3%. As a result of this increase, we also observed a corresponding decrease in the bond length between the O atoms in the polaronic octahedron and the first nearest neighbor (1NN) Ti atoms. We compute the decrease to be uniform for all such bonds, and is determined to be approximately 3.5%. As shown in Fig1(b), the in-plane bonds of tetragonal BaTiO3 are equal in length, whereas the out-of-plane bonds are of two different lengths, one short and the other long. As a polaron forms in this crystal symmetry, we observed an enhancement in the Ti-O in-plane bond lengths in the polaronic unit cell, which was computed to be 3%. However, for out-of-plane bonds, we observed that the shorter bond experienced a considerable increment of 2%, whereas a decrease of 1% was observed for the longer bond. The rhombohedral phase of BaTiO3 has two distinct bond types according to its length, three of which are longer and the other three are shorter, as shown in Fig1(c). The presence of polarons induced all three longer bonds to contract by approximately 2%, whereas all shorter bonds showed a significant elongation of approximately 8% in the polaronic site. We observed that these polaronic distortions also propagated beyond the immediate vicinity of the polaron; however, their significance gradually decreased with distance. The polaronic distortions affect the $\delta_{Fe}$, where the sites that underwent major polaronic distortions exhibited significant changes to $\delta_{Fe}$ as well. Hence, significant changes in $\delta_{Fe}$ were observed near the polaron relative to the sites farther away. Section B provides a more comprehensive discussion of the influence of electron polarons on the $\delta_{Fe}$ and properties throughout the system.
For hole polarons, the structural distortions showed slight variations compared to their electron-polaron counterparts. The bonds between the Ti atoms and polaronic O atom showed an increase in the bond length by approximately 6% for cubic BaTiO3. As a consequence of this shift of the 1NN Ti atoms, other Ti-O bonds in TiO6 octahedrons decreased, with a significant decrease of 5.5% only observed for the in-plane bond along the [100] direction. Similarly, the observations are mimicked in tetragonal BaTiO3 as a hole polaron is formed. Upon the formation of hole polarons, the length of the two bonds between the polaronic O atom and the two nearest-neighbor Ti atoms increased by approximately 4%. As a result of this increase, we observed that the other in-plane Ti-O bonds for the two neighboring octahedrons decreased in length. A significant decrease of approximately 4% is observed for one of the bonds in the [100] direction. This also results in a $\delta_{Fe}$ in the [010] direction on both 1NN TiO6 octahedrons, but the dipole moment cancels it out because they are equal and opposite. The presence of a hole polaron in the rhombohedral phase of BaTiO3 shows a similar trend; however, the elongations of the two bonds around the O atom are unequal. While one bond showed a negligible change, the other bond experienced a significant increase in length. Similar to the case with electron polarons, polaronic distortion is also observed at sites farther from the polaron, but with lesser significance relative to the nearest neighboring sites. All these polaronic distortions affect $\delta_{Fe}$ as mentioned before and this correlated effect of hole polarons on the ferroelectricity throughout the system is further discussed in section C.
III.2 Effect of electron polarons on the ferroelectric properties
The previous section discussed the influence of ferroelectricity on the formation and characteristics of the polarons. In the current section, the focus is on the reciprocal effect, that is, the impact of electron polarons on the ferroelectric properties of the tetragonal and rhombohedral phases of BaTiO3.
In tetragonal BaTiO3, the Ti atoms exist in a distorted octahedral environment with C4v symmetry. As discussed in the previous section, the self-trapping of an extra electron at the Ti site leads to significant changes in the local structure, primarily affecting the polaronic site. The electron localizes in the dxy orbital, increasing the electron density in the x-y plane and causing stronger electron-electron repulsion with the in-plane O atoms. The in-plane Ti-O bonds elongate as a result of this repulsion. Concurrently, the redistribution of electron density leads to reduced hybridization between the polaronic Ti atom and the out-of-plane O atoms, which weakens the out-of-plane Ti-O bonds. As a result, we observed that the polaronic Ti atom shifted towards the center of the octahedron by 0.014 Å, whereas all the in-plane O atoms shifted away from the Ti atom by approximately 0.07 Å. The distortions are schematically represented in the Fig.6(a). There was also a noticeable shift in both the out-of-plane O atoms by approximately 0.02 Å in the [001] direction. All these shifts result in changes in the bond lengths, as discussed Section A. The weakened hybridization between the Ti-3d and O-2p orbitals reduces the covalent character of the Ti-O bonds, decreasing the $\delta_{Fe}$ in the polaronic octahedron by approximately 15% and pushing the unit cell towards a more cubic-like shape.
For the other TiO6 octahedrons in the x-y plane containing the polaron, we observed a modest increase in the $\delta_{Fe}$, with an increase of approximately 2% for the 1NN TiO6 octahedrons in the [100], [$\bar{1}00$], [010] and [$00\bar{1}$ 0] directions. This can be attributed to the shift of in-plane oxygen away from the polaronic Ti atom. The neighboring octahedrons accommodated such shifts by shrinking one of their in-plane bonds (the Ti-O bond connecting Ti and the shifted O atoms of the polaronic octahedron) by 4%. Consequently, the Ti atom in all these 1NN octahedrons shifted further away from its centrosymmetric position by 0.01 Å. Simultaneously, the out-of-plane O atoms in the same octahedrons experienced minimal shifts to stabilize the octahedrons that were distorted owing to the shrinking of the in-plane bond. For the octahedrons in the x-y plane above the polaronic octahedron, we observed a decrease in $\delta_{Fe}$. This decrease progressively became less significant as we moved farther away from the polaronic site. A significant decrease of approximately 14% was computed for the 1NN octahedron in the [001] direction, whereas for the other sites in the same plane, the decrease was approximately 5%–8%. This can be attributed to the distortion of the out-of-plane O atom of the polaronic octahedron in the [001] direction, which results in the shortening of the longer bond of the 1NN octahedron above it. During the same time, the Ti atom in the same octahedron shifted down (that is, the direction [$00\bar{1}$]) by 0.02 Å, resulting in a significant decrease in the $\delta_{Fe}$ for this particular site. Ti shift was minimal for other sites in the same plane, which accounts for the negligible changes to the $\delta_{Fe}$. Interestingly, for the sites in the x-y plane below the polaronic octahedron, we observed little to no change in $\delta_{Fe}$. This could be due to the minimal spread of polaronic distortions in this plane owing to less susceptibility to changes since it lies in the opposite direction to the [001] tetragonal distortion in the system. However, owing to the shift in the out-of-plane O atom of the polaronic octahedron in the [001] direction, we observed an elongation of the shorter out-of-plane bond of the octahedron below the polaron. During the same time, the Ti and in-plane O atoms also shifted in the [001] direction, resulting in an overall decrease of 2% in $\delta_{Fe}$ at this site. The distortion of 0.009 Å of the Ti and O atoms is minimal in the [$00\bar{1}$] 1NN octhaedron compared to that in the [001] 1NN octahedron; hence, the observed changes to the ferroelectricity of this specific site and the others in the same plane are negligible. In summary, the variations in the $\delta_{Fe}$ exhibited planar dependency, where we clearly observed a modest increase in the x-y plane containing polarons, except for the polaronic site. For sites above this plane, there was a noticeable decrease in the $\delta_{Fe}$, whereas for sites below this plane, there was a negligible decrease in the $\delta_{Fe}$. All these conclusions are based on the results obtained with a 3x3x3 supercell.
@@ III.2 Effect of electron polarons on the ferroelectric properties
Upon the formation of electron polaron, the rhombohedral phase of BaTiO3 shows a similar trend as observed in tetragonal BaTiO3 but the reduction in $\delta_{Fe}$ is more pronounced. A substantial decrease of approximately 68% in $\delta_{Fe}$ was observed in all directions [111]. Such a decrease is the consequence of polaronic distortions, which result in a shift of the Ti atom towards its centrosymmetric position by 0.06 Å. In addition, an interesting pattern was observed for the distortions of O atoms around the polaron: the pronounced shift of O atoms in the x-direction of the octahedrons was only along the x-direction, whereas the O atoms in the y- and z-directions of the octahedron showed a similar trend of significantly shifting only along the y- and z-directions, respectively, as shown in Fig6(b). Furthermore, an asymmetric distortion, with the O atoms on either side of the polaronic Ti atom distorting non-uniformly, was observed. Specifically, a significant shift of  0.09 Å was noted for the O atoms in the [100], [010], and [001] directions, whereas a minimal shift of  0.01 Å was exhibited by the O atoms in the [$\bar{1}00$], [$00\bar{1}$ 0] and [$00\bar{1}$] directions. As there is already an existing [111] polarization in this phase of BaTiO3, the O atoms in the direction of the polarization are more prone to distortions than those in the direction opposite to the polarization. The changes in bond lengths upon polaron formation, as discussed in the Section A, can be attributed to these polaronic shifts. The Ti site holding a C3v site symmetry in rhombohedral phase creates anti-bonding states with the neighboring O-2p orbitals at the conduction band minimum. As discussed in Tsunoda et al.’s paper, electron localization on a Ti-3d orbital modifies the TiO6 octahedra such that it reduces this anti-bonding hybridization with the neighboring O-2p orbitals [60]. This observation was also similar in our calculations, and given that such reduced hybridization indicates reduced ferroelectricity for BaTiO3, this can explain the significant reduction in $\delta_{Fe}$ observed in the polaronic octahedron. The $\delta_{Fe}$ values of the neighboring sites were also affected by the polaronic distortions described earlier. For the 1NN octahedron in the [100] direction, a 0.06 Å shift of the Ti atom towards its centrosymmetric position in the x-direction was observed. This combined with the distortion of the O atom in the [100] direction of the polaronic octahedron, resulted in the shrinking of the longer bond (only in the [100] direction) by approximately 0.16 Å for this specific octahedron. Overall, a 85% decrease in $\delta_{Fe}$ in the x-direction was observed, whereas distortions along the y- and z-directions did not exhibit any variation. Furthermore, the polaronic distortions and the consequent changes to $\delta_{Fe}$ are significantly reduced with increasing distance from the polaron. The nearest neighbors in the [010] and [001] directions displayed a pattern analogous to that of the [100] nearest neighbor, with a  85% reduction in $\delta_{Fe}$ along the y- and z-directions, respectively. This can also be attributed to the polaronic distortions that resulted in the Ti atoms in the respective octahedrons shifting towards the centrosymmetric position by 0.06 Å only along the respective direction, and the directional shift of O atom in the polaronic octahedron, as described earlier. Minimal shifts of atoms present in the [$\bar{1}00$], [$00\bar{1}$ 0] and [$00\bar{1}$] directions resulted in a less pronounced reduction of 33% to $\delta_{Fe}$ across these sites.
@@
We further inspected the effect of small-polaron formation on spontaneous polarization (Ps) using the modern theory of polarization. During these calculations, we only considered the lattice effect due to polaron formation. As expected from the localized nature of the polaron-induced distortive pattern, Ps for the polaronic tetragonal BaTiO3 (0.38 C/cm2) remains close to the value of pristine BaTiO3, as tabulated in Table 1. Moreover, the energy difference between the ferroelectric and paraelectric polaronic phases was calculated as 0.032 eV/f.u., which is comparable to the energy difference of 0.03 eV/f.u. for the pristine BaTiO3 systems. Similarly, in the rhombohedral phase, the calculated Ps reduces to 0.4 C/cm2 for the polaronic state, compared to the pristine state, as tabulated in Table 1. In addition, an energy difference of 0.048 eV/f.u. in the polaronic state is comparable to 0.044 eV/f.u. in the pristine BaTiO3. The studies unveils the co-existence of $\delta_{Fe}$ and polaronic distortions, thereby revealing the formation of ferroelectric electron polarons in the material. While the self-trapping of electron polarons lead to a slight reduction in the ferroelectric properties of both the ferroelectric phases, their impact is not strong enough to completely suppress the overall ferroelectricity in the systems. The self-trapping of electron polaron is also accompanied with an introduction of local magnetism into an otherwise non-magnetic system. Magnetic moments of 0.945 $\mu{B}$ and 0.944 $\mu{B}$ were recorded for tetragonal and rhombohedral phases of BaTiO3. The reduction of normal Ti4+ to Ti3+ at polaronic site introduces an unpaired Ti-3d spin-polarized electron, resulting in magnetism that is highly concentrated at the polaronic site. Such polaron-induced magnetism in systems that are inherently ferroelectric opens up the possibility of multiferroism, which can be investigated in the future.
Table 1: Polarization values (in C/m2) for different structural phases (Cubic, Tetragonal, and Rhombohedral) of BaTiO3. The table also compares the effect of electron and hole polaron formation on the macroscopic polarization in each phase. The experimental values noted for the tetragonal and rhombohedral phases of pristine BaTiO3 are 0.26 C/m2 and 0.34 C/m2 respectively [80, 81]
Cubic Tetragonal Rhombohedral
Pristine 0 0.39 0.45
Electron Polaron 0 0.38 0.4
Hole Polaron 0 0.395 0.46
III.3 Effect of hole polarons on the ferroelectric properties
This section focuses on the impact of hole polarons on the ferroelectricity of tetragonal and rhombohedral polymorphs of BaTiO3. As observed for the case of electron polarons on tetragonal BaTiO3, the changes in the $\delta_{Fe}$ also displayed a planar dependence upon the formation of hole polarons. The tetragonal FE order was significantly affected in the octahedrons in the x-y plane containing polaronic oxygen. The modifications in the [001] $\delta_{Fe}$ were minimal for all sites in this plane, which was in contrast to from what was observed for the electron polaron described in the previous section. Owing to the shift of the two 1NN Ti atoms in the y-direction by 0.08 Å, [010] polarization is induced in both the octahedrons. However, these dipole moments nullify each other as the shifts of these Ti atoms are opposite to each other, as shown in figure 7(a), and therefore do not contribute to the net polarization of the system. In addition, as shown in figure 7(a), polaronic distortions result in the surrounding O atoms in the x-y plane moving towards polaronic oxygen, likely due to the reduced electron-electron repulsion as the electron density is reduced on the polaronic oxygen. These shifts around the polaronic O atoms are approximately 0.02 Å and are confined to the x-y plane. Consequently, we observed alterations in the FE order of the neighboring octahedrons in the x- and y-directions, introducing an additional negligible [110] distortions across these sites. However, their contribution to the overall polarization is insignificant because for each site possessing this distortion, there is an equal and opposite distortion at another site in the same plane, resulting in dipole moments canceling each other out. The changes in $\delta_{Fe}$ are tabulated in Table S3 of the supplementary material [79]. Overall, all x-y polarizations introduced to the sites in this specific plane by the hole polaron cancel out, whereas the changes in [001] distortions are negligible.
For the x-y plane above the polaron, the Ti atoms showed increased off-centering by 0.011 Å. Furthermore, the out-of-plane O atoms of the 2NN TiO6 octahedra along the y-direction show distortion towards the polaronic oxygen, resulting in an increase in the length of the longer out-of-plane bond. The combined effect is an increase of 6% in $\delta_{Fe}$ at sites belonging to this plane. Interestingly, the sites in the x-y plane below the polaron showed a negligible shift, resulting in only a minimal 3% decrease in $\delta_{Fe}$. This discrepancy in both planes could potentially be due to the pre-existing polarization direction. It is well established that the tetragonal distortion in BaTiO3 involves a downward shift of the out-of-plane O atom beneath (in the [$00\bar{1}$] direction) the Ti atom, creating a longer bond. We observed that for the octahedrons in the x-y plane above, the out-of-plane oxygen shifts inward towards the polaronic oxygen in the bottom plane, which goes hand-in-hand with the [001] tetragonal distortion.
@@ III.3 Effect of hole polarons on the ferroelectric properties
Our calculations revealed more complex alterations to $\delta_{Fe}$ upon the formation of hole polarons in the rhombohedral phase of BaTiO3. When hole polarons form on one of the in-plane O atoms of rhombohedral BaTiO3, we observed that the Ti atoms adjacent to this O atom undergo distortions, where the distortion is more pronounced on one Ti atom than on the other. In our study, the 1NN Ti atom in the [100] direction exhibited a minimal shift of 0.013 Å in the [100] direction and a 0.01 Å in the [$00\bar{1}$] direction. However, the the 1NN Ti atom in the [$\bar{1}00$] direction showed a substantial distortion of 0.17 Å along the [$\bar{1}00$] direction and a modest 0.01 Å distortion along the [$00\bar{1}$] direction. Furthermore, polaronic O atom experienced a displacement of 0.039 Å along the [100] direction. These distortions (schematically depicted in figure 7(b)) are responsible for the change in Ti-O bonds around the polaron, as discussed in section A. In addition, these distortions significantly affected the $\delta_{Fe}$ in the immediate vicinity and marginally in the farther sites. As mentioned earlier, there was a substantial shift in the [$\bar{1}00$] 1NN Ti atom, which was sufficiently pronounced to result in the reversal of the polarization in the x-direction at this site. We observed a 50% reduced negative $\delta_{Fe}$, while the $\delta_{Fe}$ also experienced a minimal 5% decrease in the z-direction for this specific site. The 1NN TiO6 octahedron in the [100] direction exhibited modest modifications in the x-direction $\delta_{Fe}$ with a 10% decrease. In addition, the y- and z-direction $\delta_{Fe}$ were modified, as tabulated in Table S4 of the supplementary material [79]. For the 2NN octahedron in the [100] direction, the Ti atom distorted by 0.05 Å towards its centrosymmetric position only along the x-direction, consequently resulting in a significant 74% reduction of the x-direction $\delta_{Fe}$. Although modest, there was also a 4% increase in the y- and z-direction $\delta_{Fe}$ for this specific octahedron.
Despite the significant changes in $\delta_{Fe}$ primarily concentrated on the octahedrons along the x-direction compared to other sites, less pronounced shifts were still observed for other sites. We observed that the 2NN along the y- and z-directions showed a clear pattern in $\delta_{Fe}$ modifications, which is correlated with the inherent [111] rhombohedral polarization. For the 2NN octahedrons along the directions aligned with a component of the [111] polarization, specifically the 2NN TiO6 octahedrons along the [010] direction aligned with the y component, and the [001] direction aligned with the z component, there was a more pronounced change in the $\delta_{Fe}$. In the [010] direction, a 17% increase in the y-direction $\delta_{Fe}$ was experienced by one site, whereas a 12% decrease was observed in the other. Similarly, both 2NN TiO6 octahedra in the [001] direction experienced a 14% decrease in the z-direction $\delta_{Fe}$. In contrast, for the 2NN along the opposite direction of the polarization component ([$00\bar{1}$ 0]] and [$00\bar{1}$]), the modifications to the $\delta_{Fe}$ were negligible. Therefore, our results indicate that inherent polarization plays a critical role in shaping the distortion pattern around the polaron, thereby affecting $\delta_{Fe}$. It appears that the atoms aligned with the direction of inherent polarization is more susceptible to transformations relative to those on the opposite side. Although the specific distortion patterns are different, these observations are similar to the case of electron polarons in rhombohedral BaTiO3.
Consistent with our approach to electron polaron systems, we also calculated Ps for hole polaron systems. Our calculations revealed a negligible effect on the overall Ps of the tetragonal phase of BaTiO3, which was determined to be 0.395 C/cm2. For the polaronic rhombohedral system, the calculations revealed a slight increase from 0.45 C/cm2 (quantified for pristine BaTiO3) to 0.46 C/cm2. Furthermore, the energy difference between the ferroelectric and paraelectric polaronic states was calculated for both phases of BaTiO3 and compared with that of pristine systems. These studies have reported a value of 0.035 eV/f.u. for tetragonal BaTiO3, which is comparable to that of the pristine states. For the rhombohedral phase, the energy difference between the polaronic states was quantified to be 0.038 eV/f.u.. Due to the formation of polarons, the hole polaron systems also revealed the presence of local magnetic moments of 0.920 $\mu{B}$, 0.926 $\mu{B}$and 0.912 $\mu{B}$ respectively for the cubic, tetragonal and rhombohedral phases.
IV Conclusion
This study looked into the interplay between ferroelectricity and polaron formation in the cubic, tetragonal and rhombohedral phases of BaTiO3. The studies reveal the stability of polaronic solution for both hole and electron polarons across all the three polymorphs of BaTiO3. Interestingly, the Ti off-centering plays a significant role in stabilizing the electron polaron formed on the Ti site for the ferroelectric phases of BaTiO3, while such an effect is not observed for hole polarons trapped on an oxygen site. Although there is only a negligible effect on overall net polarization, a slight reduction to the sponatneous polarization for the electron polaron systems and a slight increase in the hole polaron system is noticed. Since the formation of polarons is accompanied with a magnetic moment, there is a co-existence of ferroelectricity and magnetism in the systems. These findings could pave way for further optimization, potentially leading to multiferroicity in the ferroelectric phases of BaTiO3.
V ACKNOWLEDGMENTS

### v4

@@ IV.1 Polarons : Electrons and Hole Polarons in BaTiO3
Furthermore, we observed phase- and polaron-dependent modifications to the Ti-O bond lengths induced by polaron formation, providing a structural foundation for the detailed $\delta_{FE}$ analysis in Sections B-C. The formation of electron polarons generally elongates the bonds within the polaronic octahedron, consequently leading to compensatory contractions in the nearest-neighbor bonds, with distinct patterns observed for each phase. Similarly, the Ti-O bond lengths were affected by hole polaron formation, with different elongation patterns observed for different phases. The bond-length modifications for electron and hole polarons in the tetragonal and rhombohedral phases are detailed separately in Tables S1 to S4 of the SM[81]. These modifications to the bonds form the basis for the site-specific $\delta_{FE}$ changes discussed in the following sections.
IV.2 Effect of electron polarons on the ferroelectric properties
Having established the influence of ferroelectricity on polaron formation, we now examine how electron and hole polarons affect the ferroelectric properties of tetragonal and rhombohedral BaTiO3. All results below are based on calculations performed with a 3x3x3 supercell.
In tetragonal BaTiO3, the extra electron is localized in the dxy orbital of the Ti site, increasing the in-plane electron density and thereby elongating the in-plane Ti-O bonds owing to the enhanced electron-electron repulsion with the in-plane O atoms. Concurrently, the redistribution of the electron density leads to reduced hybridization between the polaronic Ti atom and out-of-plane O atoms, which weakens the out-of-plane Ti-O bonds. This results in cooperative atomic shifts: the polaronic Ti atom displaces $\sim$0.014 Å towards the octahedral center, in-plane O atoms shift outward ($\sim$0.07 Å), and out-of-plane O atoms move $\sim$0.02 Å along the [001] direction (Figure 5(a)). Consequently, these distortions result in a 15% decrease in $\delta_{FE}$ in the polaronic octahedron and push the unit cell into a cubic-like symmetry. Polaron-induced distortions propagate anisotropically to the nearest neighboring sites, with the distortion pattern influenced by the intrinsic [001] polarization of the system: The 1NN above the polaronic octahedron shows a substantial decrease in $\delta_{FE}$, while those below experience only minor suppression because the tetragonal axis hinders the distortion opposite to the polarization. This distinction aligns with the atomic relaxations depicted in Figure 5(a), where the Ti atom above the polaron shifts downward noticeably more than the Ti atom below. The 1NN octahedra in the same x-y plane showed a modest increase in $\delta_{FE}$. Table S1 of the SM details the quantitative data of the polaron-induced $\delta_{FE}$ changes[81].
Table 1: Macroscopic polarization values, relative energies, and magnetic moments for pristine, electron polaron and hole polaron states for different structural phases (Cubic, Tetragonal, and Rhombohedral) of BaTiO3. The relative energies correspond to the energy differences between the cubic and ferroelectric phases for the pristine systems and between the cubic polaron and ferroelectric polaron systems for the polaron cases. The experimental values of Ps noted for the tetragonal and rhombohedral phases of pristine BaTiO3 are 0.26 C/m2 and 0.34 C/m2 respectively [82, 83]
Phase System Spontaneous Polarization (C/m2) Relative Energy (eV/f.u.) Magnetic Moment ($\mu_{B}$)
Cubic Pristine 0.00 0.00
Electron Polaron 0.00 0.94
Hole Polaron 0.00 0.92
Tetragonal Pristine 0.39 0.030 0.00
Electron Polaron 0.38 0.032 0.95
Hole Polaron 0.395 0.035 0.926
Rhombohedral Pristine 0.45 0.044 0.00
Electron Polaron 0.4 0.048 0.95
Hole Polaron 0.46 0.038 0.912
@@ IV.2 Effect of electron polarons on the ferroelectric properties
Upon the formation of electron polaron in the rhombohedral phase of BaTiO3, pronounced anisotropic distortions are observed owing to a stronger polaron-ferroelectric coupling. The Ti site which holds a C3v site symmetry in the rhombohedral phase, creates antibonding states with the neighboring O-2pp orbitals at the conduction band minimum. As discussed in Tsunoda et al.’s study, electron localization on a Ti-3dd orbital modifies the TiO6 octahedra such that it reduces the antibonding hybridization with neighboring O-2pp orbitals[58]. This observation is similar to our calculations (in our case, polaronic Ti atom shifting towards the center of octahedron by $\sim$0.06 Å), and given that such reduced hybridization indicates reduced ferroelectricity for BaTiO3, this can explain the significant reduction of [111] $\delta_{FE}$ by 68% observed in the polaronic octahedron. The preexisting [111] polarization imposes a strong directional anisotropy on the atomic shifts, as illustrated in Figure 5(b). The O atoms (of the polaronic octahedron) in the direction of polarization ([100], [010], and [001] directions) shifted significantly by $\sim$0.09 Å, while the O atoms opposite to the direction of polarization ($\bar{1}00$], $\bar{1}0], [00$$0\bar{1}$]) displayed a comparitively minimal distortion of $\sim$0.01 Å. The polarization-biased electron-lattice coupling is reflected by this asymmetry, where the aligned O atoms experience stronger repulsive forces from the localized electron. Furthermore, the displaced O atoms within the polaronic octahedron exhibited an interesting direction-specific pattern, shifting predominantly along their directional axis. This directional anisotropy is also inherited by the neighboring octahedra, where significant changes in $\delta_{FE}$ occur at the nearest-neighbor sites aligned with the polarization direction, compared to smaller changes at sites opposite to it. Quantitative data supporting these trends are provided in Table S2 of the SM[81].
@@
This directional bias was systematically reversed when the $\delta_{FE}$ was inverted in the initial symmetry of both phases, whereas such directional anisotropy was absent in the cubic phase, confirming the role of the inherent polarization field in mediating polaronic distortions. Owing to the limited spatial range of the polaron-induced $\delta_{FE}$ pattern, the overall Ps remained near-pristine, whereas the ferroelectric-paraelectric energy differences exhibited minor shifts, as listed in Table 1. The self-trapping of electron polarons is also accompanied by the introduction of magnetism (0.94–0.95 $\mu{B}$) into an otherwise nonmagnetic system. The reduction of normal Ti4+ to Ti3+ at the polaronic site introduces an unpaired Ti-3dd spin-polarized electron, resulting in magnetism that is highly concentrated at the polaronic site. Such polaron-induced magnetism in inherently ferroelectric systems opens up the possibility of multiferroism, which can be investigated in future studies.
IV.3 Effect of hole polarons on the ferroelectric properties
For both tetragonal and rhombohedral BaTiO3, hole polaron formation leads to nuanced, but distinct, local ferroelectricity modifications. Upon local analysis of the hole polaron in the tetragonal phase, the distortions seem to reveal a planar dependence: in the x-y plane containing the polaronic O atom, the 1NN Ti atoms distort in opposite y-directions (by $\sim$0.08 Å), which induces new [010] local dipoles in these octahedra (Figure 6(a)). Because the induced dipoles are equal and opposite, they cancel each other out and do not contribute to net polarization. The surrounding O atoms in the x-y plane shift slightly (by $\sim$0.02 Å) towards the polaronic site owing to the reduced electron-electron repulsion with the polaronic O atom. These shifts, confined to the plane, also introduce additional minor [110] distortions in their respective octahedra; however, they cancel out owing to symmetry. The changes in $\delta_{FE}$ for the neighboring sites are listed in detail in Table S3 of the SM[81]. For the neighboring Ti atoms above the polaronic plane, we observed a modest increase in off-centering in the [001] direction (by $\sim$0.01 Å), resulting in a 6% increase in the local $\delta_{FE}$. However, for the Ti atoms in the plane below, the distortions were insignificant (Figure 6(a)), and the resulting changes in $\delta_{FE}$ were minimal (3% decrease). Again, such a discrepancy could be due to the influence of the pre-existing tetragonal polarization.
@@ IV.3 Effect of hole polarons on the ferroelectric properties
Our calculations revealed more complex and anisotropic alterations to $\delta_{FE}$ upon the formation of hole polarons in the rhombohedral phase of BaTiO3, as depicted in Figure 6(b). The 1NN Ti atoms undergo asymmetric distortions; one exhibits a minimal shift, while the other undergoes a substantial shift of $\sim$0.17 Å along the $0\bar{1}0$], resulting in the reversal of the local [010] polarization at that site. Furthermore, the polaronic O atom experienced displacements of $\sim$0.015 and $\sim$0.03 Å along the [010] and $\bar{1}00$] directions, respectively. These polaronic distortions resulted in significant changes in the Ti-O bonds and local $\delta_{FE}$ (detailed in Table S4 of SM) around the polaron, including a reduced 60% negative $\delta_{FE}$ (for [010] polarization) for the 1NN octahedron along the $0\bar{1}0$] direction of the polaron. The other most affected octahedron is the 2NN along the same direction, where we observed a $\sim$75% reduction (for the [010] $\delta_{FE}$ ), primarily due to the significant distortion (by $\sim$0.09 Å) of the 2NN Ti atom along the $0\bar{1}0$] direction. As observed in the case of electron polarons, the hole polaron distortions also revealed a directional dependence governed by the intrinsic polarization direction, particularly for the 2NN TiO6 octahedra in the [010] and [001] directions. The changes in these sites are clearly larger than the minimal modifications at the sites opposite to the polarization. These changes in $\delta_{FE}$ are detailed in Table S4 of the SM and are depicted schematically in Figure 6(b)[81].
Consistent with the previously described case for electron polarons, the reversal of the initial $\delta_{FE}$ naturally leads to a reversed polaronic distortion pattern, further establishing the influence of the polarization on the behavior of the polaron. The cubic phase, also for hole polarons, shows isotropic distortions because it lacks inherent polarization. Across the phases, the overall Ps and energy differences between the ferroelectric and paraelectric polaronic states remained close to the pristine values, as listed in Table 1. The hole polarons introduce local magnetic moments, as listed in Table 1, across all phases.
V Conclusion and Summary
This study investigated the interplay between ferroelectricity and polaron formation in the cubic, tetragonal, and rhombohedral phases of BaTiO3. The research involved understanding the effect of ferroelectricity on electron and hole polaron behavior in all three phases and how polarons, in turn, affect the ferroelectric properties of the systems. The findings reveal the stability of the polaronic solution for both hole and electron polarons across all thrpolaree polymorphs of BaTiO3, with the stability of electron polarons exhibiting a strong phase dependence. The polaron formation energy peaks in the rhombohedral phase due to the significant [111] Ti off-centering and the ability to reduce anti-bonding hybridisation. However, this effect is less pronounced for hole polarons trapped at oxygen sites, as they already exhibit greater stability in all phases. The study also investigated the impact of polarons on ferroelectricity, analyzing electron and hole polarons separately. It was observed that the effect on ferroelectricity is substantial at polaron localization sites and their immediate neighborhoods, whereas for sites farther from the polaron, the effect is zero or negligible. This highlights the localized nature of polarons, which significantly affect local ferroelectricity without greatly impacting macroscopic polarization. This holds true for both electron and hole polarons, although the effect on ferroelectricity in each case differs. An interesting aspect of the current work is the observed polaron-ferroelectric coupling that influences how the $\delta_{FE}$ at different sites were modified. A detailed examination revealed that the polarization field plays a significant role in polaron-induced distortions, determining the extent of modification in ferroelectric distortions due to their strong coupling. In the electron polaron tetragonal system, planar anisotropic distortions are evident, with more pronounced effects on atoms aligned with the polarization direction compared to those in the opposite direction. Similarly, in the electron polaron rhombohedral system, polaronic shifts display stronger anisotropy, with atoms aligned with each polarization component experiencing greater shifts than those in the opposite direction. Hole polarons exhibit a more complex coupling but still show anisotropy in polaronic distortions influenced by the polarization field. As polaron formation is associated with a magnetic moment, ferroelectricity and magnetism coexist in these systems. By employing the HSE06 functional to address the limitations of DFT+U, we aimed to propose a framework that, with further refinement, could potentially lead to multiferroicity in the ferroelectric phases of BaTiO3.
VI ACKNOWLEDGMENTS